The sum of two numbers is sixteen. The total of three times the smaller and twice the larger is thirty-nine.
Answer by stanbon(75887)

The sum of two numbers is sixteen. The total of three times the smaller and twice the larger is thirty-nine.
------
Equations:
x + y = 16
3x + 2y = 39
-----
3x + 3y = 3*16
3x + 2y = 39
-------
y = 9
---
x + 9 = 16
x = 7
=================
Cheers,
Stan H.
===============